III. (51 – 80) If a dumbbell has a mas of 3.63 kilograms need to be lifted,
(a) how much force is needed? If it is lifted from the ground reaching a displacement of 2.5 meters, (b) how much Work is needed?
(c) if the dumbbell is repeatedly lifted for 10 seconds, how much Power is exerted?
1
Expert's answer
2021-10-11T08:55:59-0400

(a)

"F=mg=3.63\ kg\cdot9.8\ \dfrac{m}{s^2}=35.6\ N."

(b)

"W=Fh=35.6\ N\cdot2.5\ m=89\ J."

(c)

"P=\dfrac{W}{t}=\dfrac{89\ J}{10\ s}=8.9\ W."